OpenStudy (lifeisadangerousgame):

Can someone help me with these last three calculus problems? I'm working on finding the volume of a solid that is formed by revolving the region by the x or y axis.

OpenStudy (tkhunny):

Please put enough information in the VERY FIRST post. This includes your best efforts at a solution. Many people see nothign in the first post and never come back.

OpenStudy (lifeisadangerousgame):

@tkhunny Sorry, I'll edit it

OpenStudy (arihangdu):

Uh no questions how to solve?

OpenStudy (arihangdu):

Uh no questions how to solve?

OpenStudy (tkhunny):

Sorry, disconnected... Why are you struggling with this? What have you tried? \(\pi r^{2}h\) -- The area of a circle. Differential with respect to h, \(\pi r^2\;dh\) Add them all up. \(\int\limits_{0}^{4} \pi r^{2}\;dh = \int\limits_{0}^{4} \pi (\sqrt{x})^{2}\;dx\)

OpenStudy (lifeisadangerousgame):

Oh I see. I went about it the wrong way, I had tried it in the same way you would for finding the area underneath a curve.

OpenStudy (tkhunny):

It's about the same.

OpenStudy (tkhunny):

Second one is nearly identical. Set it up!

OpenStudy (lifeisadangerousgame):

After I do (sqrtx)^2 dx, and I get just x dx, do I get the anti derivative and go from there?

OpenStudy (tkhunny):

It's an integral. How else do you plan to evaluate it? Go! You already know how to proceed. No more hesitating.

OpenStudy (tkhunny):

Third one is ALMOST the same. Just subtract the missing piece next to the y-axis. \(\pi 1^{2} 4 = 4\pi\). You'll be done in no time.

OpenStudy (lifeisadangerousgame):

8pi for the first one?

OpenStudy (tkhunny):

Not a clue. I didn't work the problem. Are you confident?

OpenStudy (lifeisadangerousgame):

Not in the least

OpenStudy (lifeisadangerousgame):

But I guess it don't really mind by now, I just want to finish everything

OpenStudy (tkhunny):

Not good enough. If you do nothing incorrectly on purpose, why would it not be correct?

OpenStudy (lifeisadangerousgame):

That's true. It should be correct then

OpenStudy (arihangdu):

👍🏻 job is done.